# Bounded Rationality Equilibrium Learning in Mean Field Games

### v1

@@ Abstract
Mean field games (MFGs) tractably model behavior in large agent populations. The literature on learning MFG equilibria typically focuses on finding Nash equilibria (NE), which assume perfectly rational agents and are hence implausible in many realistic situations. To overcome these limitations, we incorporate bounded rationality into MFGs by leveraging the well-known concept of quantal response equilibria (QRE). Two novel types of MFG QRE enable the modeling of large agent populations where individuals only noisily estimate the true objective. We also introduce a second source of bounded rationality to MFGs by restricting agents’ planning horizon. The resulting novel receding horizon (RH) MFGs are combined with QRE and existing approaches to model different aspects of bounded rationality in MFGs. We formally define MFG QRE and RH MFGs and compare them to existing equilibrium concepts such as entropy-regularized NE. Subsequently, we design generalized fixed point iteration and fictitious play algorithms to learn QRE and RH equilibria. After a theoretical analysis, we give different examples to evaluate the capabilities of our learning algorithms and outline practical differences between the equilibrium concepts.
Machine Learning, Mean Field Games
1 Introduction
@@
Learning equilibria in multi-agent games is of great practical interest but hard to scale to many agents (Daskalakis et al., 2009; Deng et al., 2023). Mean field games (MFGs) allow scaling to arbitrarily many exchangeable agents at fixed complexity. MFGs are of recent interest as a tractable method to learn approximate equilibria of rational, selfish agents (Guo et al., 2019; Cui & Koeppl, 2021; Xie et al., 2021; Laurière et al., 2022; Anahtarci et al., 2023). Thus, MFGs are applied in various settings ranging from finance to engineering (Djehiche et al., 2017; Achdou et al., 2020; Carmona, 2020).
@@
A common solution concept in the realm of multi-agent learning, and consequently in the context of MFGs, is the Nash equilibrium (NE). In a NE, each player’s strategy is considered optimal, given the strategies of the other agents, resulting in an equilibrium where no agent has an incentive to change their strategies. The optimality notion inherent in NE assumes full rationality of the individual agents.
@@
However, in many real-world situations individuals may not behave perfectly rational due to limited information processing capabilities, psychological factors, social considerations or other factors. Deviations from perfect rationality are described by the fundamental concept of bounded rationality (Simon, 1955, 1979; Kahneman & Tversky, 1982; Selten, 1990; Gigerenzer & Selten, 2002; Kahneman, 2013). Bounded rationality implies that for many real-world scenarios NE are insufficient due to their rigorous perfect rationality assumption. Instead of NE, we require a more realistic equilibrium concept accounting for partially irrational agents.
@@
A popular game-theoretic approach to modeling bounded rationality of agents are quantal response equilibria (QRE) (McKelvey & Palfrey, 1995, 1998) which are used, e.g. in economics (Breitmoser et al., 2010), robust RL (Reddi et al., 2024) and for efficient NE approximation (Gemp et al., 2024). Intuitively, in a QRE agents perceive rewards perturbed by noise and act optimally with respect to these perturbed rewards. In our work, we extend QRE to the domain of MFGs to model the behavior of a large number of agents who deviate from perfect rationality.
@@
Meanwhile on the control-theoretic side, a common approximately optimal control method is model predictive control (MPC) (Kouvaritakis & Cannon, 2016), also known as receding horizon control. To further enhance modeling of bounded rationality in MFGs, we incorporate a receding horizon method, where agents make decisions based on a limited future time horizon, reflecting more realistic decision-making processes. In contrast to MPC-based variants of MFGs such as (Inoue et al., 2021), we analyze the resulting novel receding horizon equilibria and instead focus on learning such equilibria, in a discrete-time setting.
@@
Beyond realism, introducing bounded rationality yields possible tractability advantages. NE computation for MFGs can be hard, motivating the search for alternative equilibrium notions. In this work, we show that under certain assumptions, QRE can be computed using a fixed point iteration (FPI). Moreover, QRE solutions can be seen as NE approximations with arbitrarily accurate design (Eibelshäuser & Poensgen, 2019). Recently, different equilibria have been introduced as NE approximations in MFGs (Cui & Koeppl, 2021). We compare QRE with these equilibria both theoretically and empirically and provide a new algorithm to compute QRE which extends to these equilibria. Meanwhile, for receding horizon equilibria we find novel general algorithms that work both in practice and in theory.
@@ 2 Equilibria in MFGs
2.1 Finite Agent Games
@@
2.2 Mean Field Games
MFGs are the limit of finite $N$-agent games with $N\to\infty$ and approximate many-agent finite games well. By a law of large numbers, the empirical MF is essentially replaced by its deterministic limiting MF. The idea of MFGs is to find approximate (symmetric) equilibria, which are otherwise hard to find in finite games with many agents (Deng et al., 2023). MFGs assume all agents to symmetrically play the same policy $\pi^{*}\in\Pi$ – the equilibrium solution. Whenever agent $i$ deviates from $\pi^{*}$ and instead uses some policy $\pi$, this corresponds to the policy tuple $(\pi,\underline{\pi}^{-i})$, where $\underline{\pi}^{-i}=(\pi^{*},\ldots,\pi^{*})$ denotes all but the $i$-th policy in the finite game. Hence, in the limit as $N\to\infty$ we have
@@ 2.2 Mean Field Games
2.3 Notions of Non-Cooperative Equilibria
@@
The resulting limiting NE thus becomes a policy that performs optimally against itself, i.e. when all other agents also use the same policy (Saldi et al., 2018).
@@
Proposition 1 (Saldi et al. (2018, Thm. 3.3, 4.1)).
@@
2.3.2 Regularized equilibria
A common alternative to MFNE is to use regularized control (Geist et al., 2019; Belousov & Peters, 2019) (typically entropy regularization). The idea is to replace the objective in Eq. (3) by an entropy-regularized one, which maximizes the entropy $\mathcal{H}(\hat{\pi}_{t}(\cdot\mid x_{t}))\coloneqq-\sum_{u\in\mathcal{U}}\hat{\pi}_{t}(u\mid x_{t})\log\hat{\pi}_{t}(u\mid x_{t})$ of policies in encountered states $x_{t}$,
@@ 2.3.2 Regularized equilibria
Note that for fixed $\mu\in\mathcal{M}$, it is known (e.g., Cui & Koeppl (2021)) that the optimal policy $\hat{\pi}^{\mu,\alpha}$ is
@@
There are various reasons for regularization, such as improved tractability in learning MFG equilibria (Anahtarci et al., 2023; Cui & Koeppl, 2021; Laurière et al., 2022; Li et al., 2024), exploratory properties in RL settings (Guo et al., 2022), and robustness against model uncertainty (Eysenbach & Levine, 2021).
2.3.3 Quantal response equilibria.
To incorporate bounded rationality, we assume agents act suboptimally and do not exactly optimize an objective. Building on economics literature (Breitmoser et al., 2010; Eibelshäuser & Poensgen, 2019), we introduce Markov QRE as an MFG equilibrium notion where agents only noisily estimate the state-action value function. Depending on the noise, agents act independently according to their own estimates, to the best of their knowledge. In the mentioned literature, the QRE definition is based on the state-action values $Q^{\pi}\equiv Q^{\mu,\pi}=\Gamma_{Q^{\pi}}(\mu,\pi)$ of a policy $\pi$ under current MF $\mu$, given by the recursion
@@
with $Q_{T-1}^{\pi}(x,u)=r(x,u,\mu_{T-1})$. We extend this to the optimal state-action value function $Q^{*}$ defined in Eq. (4) and denote the resulting equilibria as $\text{Q}^{\pi}\text{RE}$ and $\text{Q}^{*}\text{RE}$, respectively.
For $\text{Q}^{\pi}\text{RE}$, the noisy state-action value function is
@@
If the noise follows a Gumbel distribution with parameter $\lambda$, $\text{Q}^{\pi}\text{RE}$ and $\text{Q}^{*}\text{RE}$ policies can be computed analytically for fixed $\mu$, as softmax policies $\pi^{*}=\Gamma^{1/\lambda}_{\Pi}({Q}^{\pi})$ and $\pi^{*}=\Gamma^{1/\lambda}_{\Pi}({Q}^{*})$, respectively, where $\alpha=1/\lambda$.
The special case of $\text{Q}^{\pi}\text{RE}$ leads to the MFG analogue of the so-called logit equilibrium, considered in economics (Breitmoser et al., 2010; Eibelshäuser & Poensgen, 2019), while the special case of $\text{Q}^{*}\text{RE}$ leads to the Boltzmann equilibrium (Guo et al., 2019; Cui & Koeppl, 2021).
@@
2.3.4 Other MFG equilibrium concepts.
The literature contains many equilibrium concepts for MFGs. One example are (coarse) correlated equilibria (Campi & Fischer, 2022; Muller et al., 2022a, b) where agents obtain advice from a mediator to align their actions. The resulting correlation mechanism enables efficient equilibria calculation under moderate assumptions compared to NE. Furthermore, there are Stackelberg equilibria for MFGs (Elie et al., 2019; Carmona & Wang, 2021; Carmona et al., 2022; Vasal & Berry, 2022) where one principal tries to optimally incentivize a mean field of infinitely many agents. Since a detailed discussion of these and many more MFG equilibrium concepts is beyond the scope of our paper, we focus on NE, $\text{Q}^{\pi}\text{RE}$, $\text{Q}^{*}\text{RE}$, and RE instead and refer to the above references and Fudenberg & Tirole (1991) for a general, not MFG specific overview of equilibrium concepts.
2.3.5 Receding horizon equilibria.
Next, we describe a second method to model bounded rationality that can be combined with the previously defined equilibrium concepts. We introduce receding horizon (RH) equilibria to model limited lookahead capacity of agents by considering a shorter horizon for the underlying Bellman recursion. They describe the behaviour of agents with a model predictive control (MPC) (Kouvaritakis & Cannon, 2016), where decisions are based on a shortened future horizon, and therefore allows for more realistic or practical MFG models.
@@
where the initial MF for each MFG is the MF of the previous MFG after one time step, i.e. $\mu_{t,t}^{H}=\mu_{t-1,t}^{H}$ for all $t>0$ and $\mu_{0,0}^{H}=\mu_{0}$. $(J_{t}^{H})_{t\in\mathcal{T}}$ is the corresponding RH objective ensemble. Since MFs may deviate in practice and the horizon moves forward by one after each time step $t$, agents only implement the first entry $\pi^{*,H}_{t,t}$ of each policy. Thus, the implemented equilbrium policy $\pi^{**,H}\in\Pi$ results from the policy ensemble $(\pi^{*,H}_{t})_{t\in\mathcal{T}}\in\Pi^{T}$ by taking
@@
for each $t\in\mathcal{T}$, i.e. the diagonal $\pi^{**,H}=\mathrm{diag}((\pi^{*,H}_{t})_{t\in\mathcal{T}})$.
@@
2.4 Connections between Equilibrium Notions.
@@
In general, the equilibrium notions are distinct. Here, we look at the special cases of entropy-regularized RE and $\text{Q}^{\pi}\text{RE}$ / $\text{Q}^{*}\text{RE}$ with Gumbel noise (LQRE / BE) where the difference between $\text{Q}^{\pi}\text{RE}$, $\text{Q}^{*}\text{RE}$ and RE is the usage of policy, optimal and soft state-action value functions respectively. As $\alpha\to 0$, RE are known to essentially become NE in the sense of exploitability in the finite system (Cui & Koeppl, 2021). Furthermore, as $1/\lambda\to 0$, the Gumbel noise vanishes in $\text{Q}^{\pi}\text{RE}$ / $\text{Q}^{*}\text{RE}$. Thus, in the low temperature limit, all equilibrium notions are equivalent which is visualized in Figure 1. On the other hand, as $\alpha=1/\lambda\to\infty$, by Assm. 1, each solution tends to the uniform policy. However, for intermediate temperatures $\alpha$, the equilibria are distinct, see Figure 1. Analogously, for RH equilibria we have convergence to standard non-RH equilibria as the horizon becomes large, $H\to\infty$.
@@
3 Learning Non-Cooperative MF Equilibria
There is a variety of methods for computing or learning NE in MFGs, each with its own limitations. Standard fixed-point iteration (FPI) is not guaranteed to converge, as it cannot be Lipschitz even in simple standard finite MFGs (Cui & Koeppl, 2021, Thm. 2). Meanwhile, other algorithms such as fictitious play (FP) and Online Mirror Descent require monotonicity and their theory is currently limited to dynamics independent of the mean field (Perrin et al., 2020; Pérolat et al., 2022). Other recent ideas include an optimization-based approach (Guo et al., 2024) and a regularization approach (Guo et al., 2019; Cui & Koeppl, 2021), for which convergence via FPI is guaranteed for strong enough regularization. In the following, we generalize FPI and FP to our equilibrium concepts of interest.
@@ 3 Learning Non-Cooperative MF Equilibria
Such a result is known (Cui & Koeppl, 2021) for RE and BE ($\text{Q}^{*}\text{RE}$) and extended towards $\text{Q}^{\pi}\text{RE}$ here. As a result, the convergence of FPI follows for high temperatures.
@@
One such algorithm for standard NE is the FP algorithm. Parallel to GFPI, we formulate the generalized FP (GFP) algorithm in Alg. 2 for learning MFG equilibria. For NE, the algorithm matches with the proven FP algorithm in Perrin et al. (2020), while for $\mathrm{BE}$, RE the algorithm matches with initial experiments in Cui & Koeppl (2021). Thus, GFP is known to converge for NE under certain assumptions (Perrin et al., 2020). We extend the FP convergence proof in standard MFGs towards RE, and also towards RH equilibria by proposing suitable modifications, see the red plane in Fig. 1. As a side result, we make the existing proof for FP convergence more precise by explicitly verifying the usage of the envelope theorem, which is strictly speaking only applicable for any non-zero regularization, $\alpha>0$. The proofs for the remaining cases are left to future work. The proof requires a standard monotonicity assumption on the considered MFG, as well as a common simplifying assumption on the system dynamics and is based on a continuous-time ODE version of the algorithm with time-continuous iterations as in Perrin et al. (2020). Then, the actual algorithm can be interpreted as a discretization of the continuous-time version.
Assumption 3 (Lasry & Lions (2007)).
@@
4 Experiments
In this section we evaluate our algorithms and analyze the different equilibria for several MFGs. We analyze the efficacy of our methods for a Susceptible-Infectious-Susceptible (SIS) problem and a sequential version of Rock-Paper-Scissor (RPS) game. Additionally, we evaluate random MFGs, similar to Pérolat et al. (2022), by creating random transition and reward matrices and adding a mean-field dependent function to the reward that promotes swarm avoiding behaviour. Detailed game descriptions are found in the appendix.
@@ 4 Experiments
5 Conclusion
In this work, we introduced both QRE MFGs and RH MFGs, to incorporate bounded rationality in MFGs for a more realistic modeling of large agent populations. We compared these new equilibrium concepts to existing ones theoretically and empirically. Our analysis highlights the similarities and differences of the discussed equilibrium concepts. Furthermore, we designed general learning algorithms to compute equilibria efficiently and evaluated these algorithms on different problem settings. We hope that the novel RH and QRE MFGs combined with our learning algorithms help bring the theory closer to real-world scenarios with not perfectly rational agents. For future work, one could apply our theory and learning approach to research problems where bounded rationality is crucial, e.g. in economics or the social sciences.

### v2

@@ Abstract
Mean field games (MFGs) tractably model behavior in large agent populations. The literature on learning MFG equilibria typically focuses on finding Nash equilibria (NE), which assume perfectly rational agents and are hence implausible in many realistic situations. To overcome these limitations, we incorporate bounded rationality into MFGs by leveraging the well-known concept of quantal response equilibria (QRE). Two novel types of MFG QRE enable the modeling of large agent populations where individuals only noisily estimate the true objective. We also introduce a second source of bounded rationality to MFGs by restricting the agents’ planning horizon. The resulting novel receding horizon (RH) MFGs are combined with QRE and existing approaches to model different aspects of bounded rationality in MFGs. We formally define MFG QRE and RH MFGs and compare them to existing equilibrium concepts such as entropy-regularized NE. Subsequently, we design generalized fixed-point iteration and fictitious play algorithms to learn QRE and RH equilibria. After a theoretical analysis, we give different examples to evaluate the capabilities of our learning algorithms and outline practical differences between the equilibrium concepts.
Introduction
@@
Learning equilibria in multi-agent games is of great practical interest but hard to scale to many agents (Daskalakis, Goldberg, and Papadimitriou 2009; Deng et al. 2023). Mean field games (MFGs) allow scaling to arbitrarily many exchangeable agents at fixed complexity. MFGs are of recent interest as a tractable method to learn approximate equilibria of rational, selfish agents (Guo et al. 2019; Cui and Koeppl 2021; Xie et al. 2021; Laurière et al. 2022; Anahtarci, Kariksiz, and Saldi 2023). Thus, MFGs are applied in various settings ranging from finance to engineering (Djehiche, Tcheukam, and Tembine 2017; Achdou et al. 2020; Carmona 2020).
@@
A common concept in multi-agent learning is the Nash equilibrium (NE), where each player’s strategy is optimal given others’, leading to no incentive for agents to change strategies. The optimality notion inherent in NE assumes full rationality of the individual agents.
@@
However, in many real-world situations individuals may not behave perfectly rational due to limited information processing capabilities, psychological factors, social considerations or other factors. Deviations from perfect rationality are described by the fundamental concept of bounded rationality (Simon 1955, 1979; Kahneman and Tversky 1982; Selten 1990; Gigerenzer and Selten 2002; Kahneman 2013). Bounded rationality implies that for many real-world scenarios NE are insufficient due to their rigorous perfect rationality assumption. Instead of NE, we require a more realistic equilibrium concept accounting for partially irrational agents.
@@
A popular game-theoretic approach to modeling bounded rationality of agents are quantal response equilibria (QRE) (McKelvey and Palfrey 1995, 1998) which are used, e.g., in economics (Breitmoser, Tan, and Zizzo 2010), robust RL (Reddi et al. 2024) and for efficient NE approximation (Gemp, Marris, and Piliouras 2024). Intuitively, in a QRE agents perceive rewards perturbed by noise and act optimally with respect to these perturbed rewards. In our work, we extend QRE to the domain of MFGs to model the behavior of a large number of agents who deviate from perfect rationality.
@@
Meanwhile, on the control-theoretic side, a common approximately optimal control method is model predictive control (MPC) (Kouvaritakis and Cannon 2016), also known as receding horizon control. To further enhance modeling of bounded rationality in MFGs, we incorporate a receding horizon method, where agents make decisions based on a limited future time horizon, reflecting more realistic decision-making processes. In contrast to MPC-based variants of MFGs such as (Inoue et al. 2021), we analyze the resulting novel receding horizon equilibria and instead focus on learning such equilibria, in a discrete-time setting.
@@
Beyond realism, introducing bounded rationality yields possible tractability advantages. NE computation for MFGs can be hard, motivating the search for alternative equilibrium notions. We show that under certain assumptions, QRE can be computed using a fixed-point iteration (FPI). Moreover, QRE solutions can be seen as NE approximations with arbitrarily accurate design (Eibelshäuser and Poensgen 2019). Recently, different equilibria have been introduced as NE approximations in MFGs (Cui and Koeppl 2021). We compare QRE with these equilibria theoretically and empirically and provide a new algorithm to compute QRE which extends to these equilibria. For receding horizon equilibria, we develop novel algorithms effective in theory and practice.
@@
Finite Agent Games
@@
Mean Field Games
MFGs are the limit of finite $N$-agent games with $N\to\infty$ and approximate many-agent finite games well. By a law of large numbers, the empirical MF is essentially replaced by its deterministic limiting MF. The idea of MFGs is to find approximate (symmetric) equilibria, which are otherwise hard to find in finite games with many agents (Deng et al. 2023). MFGs assume all agents to symmetrically play the same policy $\pi^{*}\in\Pi$ – the equilibrium solution. Whenever agent $i$ deviates from $\pi^{*}$ and instead uses some policy $\pi$, this corresponds to the policy tuple $(\pi,\underline{\pi}^{-i})$, where $\underline{\pi}^{-i}=(\pi^{*},\ldots,\pi^{*})$ denotes all but the $i$-th policy in the finite game. Hence, in the limit as $N\to\infty$ we have
@@
Notions of Non-Cooperative Equilibria
@@
The resulting limiting NE thus becomes a policy that performs optimally against itself, i.e. when all other agents also use the same policy (Saldi, Basar, and Raginsky 2018).
@@
Proposition 1 (Saldi, Basar, and Raginsky (2018, Thm. 3.3, 4.1)).
@@
Regularized equilibria.
A common alternative to MFNE is to use regularized control (Geist, Scherrer, and Pietquin 2019; Belousov and Peters 2019) (typically entropy regularization). The idea is to replace the objective in Eq. (3) by an entropy-regularized one, which maximizes the entropy $\mathcal{H}(\hat{\pi}_{t}(\cdot\mid x_{t}))\coloneqq-\sum_{u\in\mathcal{U}}\hat{\pi}_{t}(u\mid x_{t})\log\hat{\pi}_{t}(u\mid x_{t})$ of policies in encountered states $x_{t}$,
@@
Note that for fixed $\mu\in\mathcal{M}$, it is known (e.g., Cui and Koeppl (2021)) that the optimal policy $\hat{\pi}^{\mu,\alpha}$ is
@@
There are various reasons for regularization, such as improved tractability in learning MFG equilibria (Anahtarci, Kariksiz, and Saldi 2023; Cui and Koeppl 2021; Laurière et al. 2022; Li et al. 2024), exploratory properties in RL settings (Guo, Xu, and Zariphopoulou 2022), and robustness against model uncertainty (Eysenbach and Levine 2021).
Quantal response equilibria.
To incorporate bounded rationality, we assume agents act suboptimally and do not exactly optimize an objective. Building on economics literature (Breitmoser, Tan, and Zizzo 2010; Eibelshäuser and Poensgen 2019), we introduce Markov QRE as an MFG equilibrium notion where agents only noisily estimate the state-action value function. Depending on the noise, agents act independently according to their own estimates, to the best of their knowledge. In the mentioned literature, the QRE definition is based on the state-action values $Q^{\pi}\equiv Q^{\mu,\pi}=\Gamma_{Q^{\pi}}(\mu,\pi)$ of a policy $\pi$ under current MF $\mu$, given by the recursion
@@
with $Q_{T-1}^{\pi}(x,u)=r(x,u,\mu_{T-1})$. We extend this to the optimal state-action value function $Q^{*}$ defined in Eq. (4) and denote the resulting equilibria as $\text{Q}^{\pi}\text{RE}$ and $\text{Q}^{*}\text{RE}$, respectively. For $\text{Q}^{\pi}\text{RE}$, the noisy state-action value function is
@@
If the noise follows a Gumbel distribution with parameter $\lambda$, $\text{Q}^{\pi}\text{RE}$ and $\text{Q}^{*}\text{RE}$ policies can be computed analytically for fixed $\mu$, as softmax policies $\pi^{*}=\Gamma^{1/\lambda}_{\Pi}({Q}^{\pi})$ and $\pi^{*}=\Gamma^{1/\lambda}_{\Pi}({Q}^{*})$, respectively, where $\alpha=1/\lambda$. The special case of $\text{Q}^{\pi}\text{RE}$ leads to the MFG analogue of the so-called logit equilibrium, considered in economics (Breitmoser, Tan, and Zizzo 2010; Eibelshäuser and Poensgen 2019), while the special case of $\text{Q}^{*}\text{RE}$ leads to the Boltzmann equilibrium (Guo et al. 2019; Cui and Koeppl 2021).
@@
Other MFG equilibrium concepts.
The literature contains many equilibrium concepts for MFGs. One example are (coarse) correlated equilibria (Campi and Fischer 2022; Muller et al. 2022a, b) where agents obtain advice from a mediator to align their actions. The resulting correlation mechanism enables efficient equilibria calculation under moderate assumptions compared to NE. Furthermore, there are Stackelberg equilibria for MFGs (Elie, Mastrolia, and Possamaï 2019; Carmona and Wang 2021; Carmona, Dayanıklı, and Laurière 2022; Vasal and Berry 2022) where one principal tries to optimally incentivize a mean field of infinitely many agents. Since a detailed discussion of these and many more MFG equilibrium concepts is beyond the scope of our paper, we focus on NE, $\text{Q}^{\pi}\text{RE}$, $\text{Q}^{*}\text{RE}$, and RE instead and refer to the above references and Fudenberg and Tirole (1991) for a general, not MFG specific overview of equilibrium concepts.
Receding horizon equilibria.
Next, we describe a second method to model bounded rationality that can be combined with the previously defined equilibrium concepts. We introduce receding horizon (RH) equilibria to model limited lookahead capacity of agents by considering a shorter horizon for the underlying Bellman recursion. They describe the behaviour of agents with a model predictive control (MPC) (Kouvaritakis and Cannon 2016), where decisions are based on a shortened future horizon, and therefore allows for more realistic or practical MFG models.
@@
where the initial MF for each MFG is the MF of the previous MFG after one time step, i.e. $\mu_{t,t}^{H}=\mu_{t-1,t}^{H}$ for all $t>0$ and $\mu_{0,0}^{H}=\mu_{0}$. $(J_{t}^{H})_{t\in\mathcal{T}}$ is the corresponding RH objective ensemble. Since MFs may deviate in practice and the horizon moves forward by one after each time step $t$, agents only implement the first entry $\pi^{*,H}_{t,t}$ of each policy. Thus, the implemented equilibrium policy $\pi^{**,H}\in\Pi$ results from the policy ensemble $(\pi^{*,H}_{t})_{t\in\mathcal{T}}\in\Pi^{T}$ by taking
@@
for each $t\in\mathcal{T}$, i.e., the diagonal $\pi^{**,H}=\mathrm{diag}((\pi^{*,H}_{t})_{t\in\mathcal{T}})$.
@@
Connections between Equilibrium Notions
@@
In general, the equilibrium notions are distinct. Here, we look at the special cases of entropy-regularized RE and $\text{Q}^{\pi}\text{RE}$ / $\text{Q}^{*}\text{RE}$ with Gumbel noise (LQRE / BE) where the difference between $\text{Q}^{\pi}\text{RE}$, $\text{Q}^{*}\text{RE}$ and RE is the usage of policy, optimal and soft state-action value functions respectively. As $\alpha\to 0$, RE are known to essentially become NE in the sense of exploitability in the finite system (Cui and Koeppl 2021). Furthermore, as $1/\lambda\to 0$, the Gumbel noise vanishes in $\text{Q}^{\pi}\text{RE}$ / $\text{Q}^{*}\text{RE}$. Thus, in the low temperature limit, all equilibrium notions are equivalent which is visualized in Figure 1. On the other hand, as $\alpha=1/\lambda\to\infty$, by Assm. 1, each solution tends to the uniform policy. However, for intermediate temperatures $\alpha$, the equilibria are distinct, see Figure 1. Analogously, for RH equilibria we have convergence to standard non-RH equilibria as the horizon becomes large, $H\to\infty$.
@@
Learning Non-Cooperative MF Equilibria
There is a variety of methods for computing or learning NE in MFGs, each with its own limitations. Standard fixed-point iteration (FPI) is not guaranteed to converge, as it cannot be Lipschitz even in simple standard finite MFGs (Cui and Koeppl 2021, Thm. 2). Meanwhile, other algorithms such as fictitious play (FP) and Online Mirror Descent require monotonicity and their theory is currently limited to dynamics independent of the mean field (Perrin et al. 2020; Pérolat et al. 2022). Other recent ideas include an optimization-based approach (Guo, Hu, and Zhang 2024) and a regularization approach (Guo et al. 2019; Cui and Koeppl 2021), for which convergence via FPI is guaranteed for strong enough regularization. In the following, we generalize FPI and FP to our equilibrium concepts of interest.
@@
Such a result is known (Cui and Koeppl 2021) for RE and BE ($\text{Q}^{*}\text{RE}$) and extended towards $\text{Q}^{\pi}\text{RE}$ here. As a result, the convergence of FPI follows for high temperatures.
@@
One such algorithm for standard NE is the FP algorithm. Parallel to GFPI, we formulate the generalized FP (GFP) algorithm in Alg. 2 for learning MFG equilibria. For NE, the algorithm matches with the proven FP algorithm in Perrin et al. (2020), while for $\mathrm{BE}$, RE the algorithm matches with initial experiments in Cui and Koeppl (2021). Thus, GFP is known to converge for NE under certain assumptions (Perrin et al. 2020). We extend the FP convergence proof in standard MFGs towards RE, and also towards RH equilibria by proposing suitable modifications, see the red plane in Fig. 1. As a side result, we make the existing proof for FP convergence more precise by explicitly verifying the usage of the envelope theorem, which is strictly speaking only applicable for any non-zero regularization, $\alpha>0$. The proofs for the remaining cases are left to future work. The proof requires a standard monotonicity assumption on the considered MFG, as well as a common simplifying assumption on the system dynamics and is based on a continuous-time ODE version of the algorithm with time-continuous iterations as in Perrin et al. (2020). Then, the actual algorithm can be interpreted as a discretization of the continuous-time version.
Assumption 3 (Lasry and Lions (2007)).
@@
Experiments
In this section we evaluate our algorithms and analyze the different equilibria for several MFGs. We analyze the efficacy of our methods for a Susceptible-Infectious-Susceptible (SIS) problem and a sequential version of Rock-Paper-Scissor (RPS) game. Additionally, we evaluate random MFGs, similar to Pérolat et al. (2022), by creating random transition and reward matrices and adding a mean-field dependent function to the reward that promotes swarm avoiding behaviour. Detailed game descriptions are found in the appendix. For code, see https://github.com/yannickeich/QRE-MFG.
@@
Conclusion
In this work, we introduced both QRE MFGs and RH MFGs, to incorporate bounded rationality in MFGs for a more realistic modeling of large agent populations. We compared these new equilbrium concepts to existing ones theoretically and empirically. Our analysis highlights the similarities and differences of the discussed equilibrium concepts. Furthermore, we designed general learning algorithms to compute equilibria efficiently and evaluated these algorithms on different problem settings. We hope that the novel RH and QRE MFGs combined with our learning algorithms help bring the theory closer to real-world scenarios with not perfectly rational agents. For future work, one could apply our theory and learning approach to research problems where bounded rationality is crucial, e.g. in economics or the social sciences.